Circular No. 5619

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992ay IN ANONYMOUS GALAXY J. Maza, University of Chile; and M. Hamuy, Cerro Tololo Interameri-

can Observatory, report the discovery by R. Antezana (University of Chile) of an apparent supernova located in a system of two anonymous galaxies (α = 4^{h}28^{m}50^{s}.2, δ = −46^{o}30′32′′, equinox 1950.0). Its position is 6′′ east and 6′′ north of the southernmost galaxy. The supernova, of m_{pg} ∼ 18.5, was found on a 20-min unfiltered IIa-O plate taken by C. Muena (CTIO) with the CTIO Curtis Schmidt telescope on Sept. 20.323 UT. Confirmation was made by X. Guo (Yale University) on CCD B and V images obtained on Sept. 25.403 with the CTIO 0.9-m telescope.

GRS 1915+105 B. A. Harmon, W. S. Paciesas, and G. J. Fishman, report for the

BATSE/Compton Observatory Team: “GRS 1915+105, the hard x-ray transient discovered by the Granat WATCH experiment (IAUC 5590), has been observed by BATSE, at flux levels between 280 and 400 mCrab since Aug. 15 in the energy band 20-230 keV, and continues to be in outburst. A flux history derived for the given location indicates the following levels: June 16-24, 90 ± 10 mCrab; July 10-16, 190 ± 20 mCrab; July 17-Aug. 5, 280 ± 30 mCrab. Systematic effects due to source confusion appear minimal in these time intervals. The source has remained at ∼ 300 mCrab since mid-August. Power-law fits to the spectrum in the band 20-120 keV from June to September indicate an evolution from spectral index –2.0 (June) to –2.8 (September). The spectral characteristics and slow rise to apparent maximum intensity are similar to that observed during the 1991 outburst of the black-hole candidate GX 339-4 (IAUC 5327, 5352), but differ markedly from those of the other recent transient GRO J0422+32 (IAUC 5580, 5584).”

FG SAGITTAE C. E. Woodward, J. Cole, and T. Hodge, University of Wyoming; and

R. D. Gehrz, University of Minnesota, write: “We have started to monitor the decline in the light curve of FG Sge (IAUC 5610) and report the following infrared magnitudes, obtained on the 2.3-m Wyoming Infrared Observatory telescope: Sept. 17.22 UT (with bolometer detector), M = 3.64 ± 0.10, N = 2.91 ± 0.16; Sept. 19.14 (with InSb detector), J = 8.69 ± 0.12, H = 7.69 ± 0.05, K = 6.26 ± 0.05, L′ = 4.08 ± 0.01.”

1992 September 26 Daniel W. E. Green